On January 1, Park Corporation and Strand Corporation had condensed balance sheets as follows:

Park Strand
Current assets $ 92,500 $ 31,050
Noncurrent assets 114,500 49,700
Total assets $ 207,000 $ 80,750
Current liabilities $ 44,000 $ 30,750
Long-term debt 64,000
Stockholders' equity 99,000 50,000
Total liabilities and equities $ 207,000 $ 80,750

On January 2, Park borrowed $58,400 and used the proceeds to obtain 80 percent of the outstanding common shares of Strand. The acquisition price was considered proportionate to Strand’s total fair value. The $58,400 debt is payable in 10 equal annual principal payments, plus interest, beginning December 31. The excess fair value of the investment over the underlying book value of the acquired net assets is allocated to inventory (60 percent) and to goodwill (40 percent).

1) On a consolidated balance sheet as of January 2, what should be the amount for current assets?

2) On a consolidated balance sheet as of January 2, what should be the amount for noncurrent assets?

3) On a consolidated balance sheet as of January 2, what should be the amount for current liabilities?

4) On a consolidated balance sheet as of January 2, what should be the amount for noncurrent liabilities?

5)On a consolidated balance sheet as of January 2, what should be the amount for stockholders' equity?
Answer #1

Please find below table useful to compute desired results: -

B i Particulars Amount 2 Fair Value on Acquisition date =58400/0.8 3 Less: Book Value of S =31050+49700-30750 4 Fair Value in

End results would be as follows: -

A в 1 Particulars Amount 2 Fair Value on Acquisition date $73,000.00 3 Less: Book Value of S $50,000.00 4 Fair Value in Exces